OpenStudy (anonymous):

A rectangular plot of land will be bounded on one side by a river. On the other three sides, it will be bounded by a fence. With 500m of fence to use, what is the largest area you can enclose, and what are its dimensions? (I'm having trouble with this one)

OpenStudy (anonymous):

I do know the more closely it becomes a square the larger the area you have. Let me think on this and get back to you shortly.

OpenStudy (anonymous):

A = Length * width A = lw 500 = 2w + l solving this for l you get l = 500 - 2w substitute that back into the A=lw A = (500 - 2w)w distribute A = 500w - 2w^2 The maximum can be found at the vertex which you find the w value by -b/2a = -500/-2(2) = w = 125 so I think the width = 125 and the length would be 250

OpenStudy (amistre64):

2x + y = 500 y = 500-2x area = xy A = x(500-2x)

OpenStudy (anonymous):

I would find the maximum by completing the square, but I'm a badass.

OpenStudy (amistre64):

A = 500x -2x^2 dA = 500 -4x x = 500/4 = 250/2 = 125

OpenStudy (anonymous):

Thank you guys for your help...You are all awesome =D

OpenStudy (amistre64):

y = 500-2x y = 500 - 2(125) y = 500 - 250 y = 250

OpenStudy (amistre64):

blext was right...and first :)

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

But you did it the "cool" way